OpenStudy (anonymous):

Please help!!! Harrison and Sherrie are making decisions on their bank accounts. Harrison wants to put more money in as a principle amount because the more you start with, the more interest you will gain. Sherrie wants to put the original money in an account with a higher interest rate. Explain which method will result in more money.

OpenStudy (amistre64):

depends on the interest rates ...

OpenStudy (amistre64):

B(1+i+k)^n = (B+d)(1+i)^n if: B/(B+d) = (1+i)^n/(1+i+k)^n then your makeing the same exact amount of money per time frame

OpenStudy (anonymous):

They don't give any numbers in this question. It makes it that much more confusing

OpenStudy (amistre64):

numbers dont solve things, concepts do. numbers are for a specific individual instant in time; generalities cover broader scopes and situations

OpenStudy (amistre64):

you have a starting amount B, and an interest rate i, and a time frame of n periods B(1+i)^n is how much you are making at the moment, so modify it B+d = more deposit i+k = more interest

OpenStudy (anonymous):

Okay...

OpenStudy (amistre64):

when are they the same? when they equal each other of course more deposit = higher interest (B+d)(1+i)^n = B(1+i+k)^n then its basic math from there

OpenStudy (amistre64):

so it depends in the amount deposited, and the interests involved. a case can be made either way depending in what the now specific situation provides

OpenStudy (anonymous):

Would Harrison be correct overall?

OpenStudy (amistre64):

neither of them are correct, or incorrect. But we have a way to determine it when we know their situation

OpenStudy (amistre64):

we have developed a 'formula' that can be used as proof.

OpenStudy (amistre64):

spose we start with 100 at 5% interest for 12 periods, and the guy says add 50; what interest rate does the wife need to find to win the deal?

OpenStudy (amistre64):

B/(B+d) = (1+i)^n/(1+i+k)^n 100/150 = ((1+.05)/(1.05+k))^12 well we solve and find that this is equal when k is about 3.6% if the wife finds an investment for 5+4 = 9% she wins, otherwise she loses.

OpenStudy (anonymous):

I do not understand how that is relevant to the question

OpenStudy (amistre64):

Explain which method will result in more money. i just proved that the situation can go either way ...

OpenStudy (amistre64):

the question isnt who wins ... the question is asking you to think about it and come up with a suitable process that can determine an outcome.

OpenStudy (amistre64):

critical thinking skills they call it

OpenStudy (anonymous):

Oh okay I get it. Thanks!

OpenStudy (amistre64):

good luck